OpenStudy (anonymous):

A rectangular plot of land is 100 feet long and 50 feet wide. How long is the walkway along the diagonal? Round to the nearest foot. A. 75 feet B. 87 feet C. 112 feet D. 150 feet

jimthompson5910 (jim_thompson5910):

hint:  \[\Large x^2 + y^2 = z^2\]

jimthompson5910 (jim_thompson5910):

In this case, x = 100 y = 50 z = unknown (for now)

OpenStudy (anonymous):

100^2 + 50^2 = 111.80 ?

jimthompson5910 (jim_thompson5910):

You take the square root of 100^2 + 50^2, yes (I'm sure you did that but you didn't show it)

jimthompson5910 (jim_thompson5910):

so round that to the nearest whole number to get 112